# All Optical Echo State Network Reservoir Computing

### v1

@@ 4.2 NARMA10
where $y_{k}$ is the output at time step $k$ for input $u_{k}$. We have used values for the constant parameters $(\alpha,\beta,\gamma,\delta)=(0.3,0.05,1.5,0.1)$ and $\Gamma_{k}$ is a random variable that takes values from the interval $[-1,1]$. We have used $\mu=0$ and $\kappa=0.1$. This $y_{k}$ given by Equations (9) and (10) gives us the target sequence that we fit the ESN estimated output to. This is a tenth-order equation with a large time delay of 10 steps, making it a difficult problem. It demonstrates the capability of the reservoir computer to estimate polynomials given random inputs. The results for a software-based ESN are used to evaluate the performance of the optical ESN.
@@
(a) Average NMSE vs. Pump Splitting ratio for various $gL$, $N=10$.
(b) Average NMSE vs. Pump Splitting ratio for various $N$, $gL=12$.
(c) Average NMSE vs. $gL$ values for various $N$, $\mathrm{Pump/Input}=1400$.
@@
4.3 Mackey-Glass
@@
(a) Software (Sigmoid) Mackey Glass Test, NMSE = 0.04.
(b) Optical Mackey Glass Test, NMSE = 0.06.
(a) Average NMSE vs. Pump Splitting ratio for various $gL$, $N=10$.
(b) Average NMSE vs. Pump Splitting ratio for various $N$, $gL=12$.
(c) Average NMSE vs. $gL$ values for various $N$, $\mathrm{Pump/Input}=1200$.
@@
Figure 8 shows the behavior with the variation of pump splitting ratio and $gL$ value. Similar to the results from the sine square classification test, the NMSE decreases with larger pump splitting ratio and shows improper fitting for $N=4$. We also see a similar dip in NMSE in lower $gL$ values. So, across all tests, it seems that a high pump splitting ratio and a $gL$ value of around $5\;W^{-1}$ can ensure the best results. While these behaviors can vary between tests and randomization, ensuring good performance across all three tests ensures good performance for the optical ESN in general application.
4.4 Noise Performance
For the case of a weak input probe undergoing high gain amplification, Brillouin amplification is known to be noisy [32]. This noise scales with pump intensity and is more prominent in the case of pump beam aberration, which can cause reduced performance by the optical reservoir [33]. In our design, these intense gain values are not necessary for the matrix multiplication steps, making our reservoir relatively safe from high gain noise. To further reduce the impact, the amplified spontaneous noise (ASE) can be mitigated by the use of a band-pass filter [22] and in our simulations, we see good performance without requiring output probe powers over 40 dBm, which would result in a negligible noise level.
@@ 4.4 Noise Performance
(b) Optical NARMA Test With Noise,
NMSE=0.75.
@@
(a) Average NMSE of Sine Square Classification vs. Pump Splitting ratio for various $N$, $gL=12$.
(b) Average NMSE of the NARMA Test vs. Pump Splitting ratio for various $N$, $gL=12$.
(c) Average NMSE of the Mackey-Glass Test vs. Pump Splitting ratio for various $N$, $gL=12$.
@@
Another source of noise is optical power detection noise. This is accounted for by a simple gaussian noise model. Figure 9 shows examples of the performance of the optical ESN with detector noise included, compared with Figures 3, 5, and 7. For this test, we assumed the noise property of $0.32\%$ uncertainty at $1550$nm wavelength, which is a standard treatment of considering detector noise [34]. In the presence of noise, the nonlinear capability is severely reduced as can be seen in the sine-square classification test results. This is mitigated by choosing an optimal $gL$ value with a large pump splitting ratio, as detailed in the previous section. We observe from Figure 10 that the average NMSE begins to trend upward, unlike the noise free cases of Figures 4, 6, and 8. The cause of this is likely that the saturation point for the specific task as a function of pump/input ratio is reached earlier because of the presence of detection noise.
@@
The optical ESN shows robustness in its time-series prediction ability, even in the presence of noise. The reported median NMSE in Figure 10 for the Mackey-Glass test is 0.15. This shows that the long-term memory capability is not diminished by the presence of noise, although there is a plateau in performance with varying pump splitting ratio, as can be seen from Figure 10 (c). We can see that overall, after accounting for power detection noise in the fiber setup, there is a reduction in performance of around $0.1-0.5$ NMSE with more reduction in the classification task and less effect on prediction, implying that the efficiency of nonlinearity is affected more by the noise than the long and short term memory capabilities of the optical reservoir. These noise properties are for commercially available optical detectors. Hence, the negative impact of noise on performance can be further mitigated with the use of low noise detectors [35], [36].
5 Conclusion
@@
We have provided the design for a feasible all-optical ESN with minimal resource overhead and energy requirements. For the nonlinear activation and matrix multiplication, we use SBS in two regimes: the large pump regime for linear behavior, and the self-pumped regime for nonlinear behavior. In the linear regime, tuning the pump power is enough to achieve the necessary multiplications. For the nonlinear activation step, according to reported achievable values of Brillouin gain in highly nonlinear fibers (HNLFs) [22], our design requires a fiber of length $3-6$ m to achieve the required nonlinearity for good results, which is a very practical length of fiber. Such fibers also have a reasonable threshold power to show the required nonlinear output.
@@
Additionally, our design requires only a single Brillouin fiber amplifier as a nonlinear activation device, making it highly scalable. However, a trade-off in computation speed has been introduced to allow for this scalable design. It is possible to achieve faster operation speeds at the cost of scalability, but this also greatly increases the hardware resource requirement, so we have decided against this approach. The nonlinear activation is optimizable by tuning the $gL$ value of the SBS amplifier as well as the pump splitting ratio, which refers to the ratio of the pump to the input given to the self-pumped Brillouin amplifier. By doing so, we analyze the behavior of the optical ESN for multiple benchmark tests to find the most optimal activation function design.
@@
Our matrix multiplication and beam combination steps make use of standard fiber optical components such as VOAs, fiber beam combiners, 1:N splitters, combiners, and a 1:N switch. We see comparable results to a standard software ESN using a sigmoid activation function. To show this, we chose three tasks: classification of sine and square waves, prediction of the NARMA10 sequence, and prediction of the Mackey-Glass chaotic time series. Each benchmark task demonstrates the nonlinearity, polynomial algebra capability, and long-term memory of this ESN, respectively, showing that universal computations are indeed accessible with this hardware. We validated our numerical simulations of the designed all-optical ESNs by comparing them to a software ESN with a sigmoid activation function. The results demonstrated that both systems achieve similar performance when using an equal number of nodes. This shows the feasibility and effectiveness of an all-optical approach to building an ESN.
[…]
Disclosures The authors declare no conflicts of interest.
[…]
Data availability Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
@@ Appendix A SBS Transcendental Equation Derivation
[…]
$\implies ln\left(\frac{P_{\mathrm{s}}(z)(P_{\mathrm{s}}(z)+P_{\mathrm{p}}-P_{\mathrm{out}})}{P_{\mathrm{in}}(P_{\mathrm{s}}(z)+P_{\mathrm{p}}-P_{\mathrm{out}})}\right)=-gCz.$ (16)
@@
$=(P_{\mathrm{s}}(z)P_{\mathrm{in}}-P_{\mathrm{out}}P_{\mathrm{in}})e^{-(P_{\mathrm{p}}-P_{\mathrm{out}})gz}.$ (17)

### v2

@@ IV.2 NARMA10
where $y_{k}$ is the output at time step $k$ for input $u_{k}$. We have used values for the constant parameters $(\alpha,\beta,\gamma,\delta)=(0.3,0.05,1.5,0.1)$ and $\Gamma_{k}$ is a random variable that takes values from the interval [1,1][-1,1]. We have used $\mu=0$ and $\kappa=0.1$. This $y_{k}$ given by Equations (9) and (10) gives us the target sequence that we fit the ESN estimated output to. This is a tenth-order equation with a large time delay of 10 steps, making it a difficult problem. It demonstrates the capability of the reservoir computer to estimate polynomials given random inputs. The results for a software-based ESN are used to evaluate the performance of the optical ESN.
@@
(a) Software (Sigmoid) Mackey Glass Test, NMSE = 0.04.
(b) Optical Mackey Glass Test, NMSE = 0.06.
IV.3 Mackey-Glass
@@
(a) Average NMSE vs. Pump Splitting ratio for various gLgL, $N=10$.
(b) Average NMSE vs. Pump Splitting ratio for various NN, $gL=12$.
(c) Average NMSE vs. gLgL values for various NN, $\mathrm{Pump/Input}=1000$.
@@
(b) Optical NARMA Test With Noise, NMSE=0.75.
@@
(a) Average NMSE of Sine Square Classification vs. Pump Splitting ratio for various NN, $gL=12$.
(b) Average NMSE of the NARMA Test vs. Pump Splitting ratio for various NN, $gL=12$.
(c) Average NMSE of the Mackey-Glass Test vs. Pump Splitting ratio for various NN, $gL=12$.
@@
Figure 8 shows the behavior with the variation of pump splitting ratio and $gL$ value. Similar to the results from the sine square classification test, the NMSE decreases with larger pump splitting ratio and shows improper fitting for $N=4$. Although, the variation in values is slight here as well, similar to the NARMA10 test. Across all tests, it seems that a high pump splitting ratio and a $gL$ value of around $5-10\;W^{-1}$ can ensure the best results, as the improvement from further increasing the $gL$ value is marginal at the cost of increased hardware complexity. While these behaviors can vary between tests and randomization, ensuring good performance across all three tests ensures good performance for the optical ESN in general application. In order to limit the energy consumption of the system, the pump splitting ratio can be limited to 1000. This still shows good results, as can be seen from Figures 4, 6, and 8. While a larger pump splitting ratio is generally more beneficial to performance, we accept the slight degradation in exchange for enhanced feasibility of the system through a smaller pump power budget.
IV.4 Noise Performance and Negative Values
For the case of a weak input probe undergoing high gain amplification, Brillouin amplification is known to be noisy [36]. This noise scales with pump intensity and is more prominent in cases with pump beam aberration, which can cause reduced performance by the optical reservoir [37]. In our design, these intense gain values are not necessary for the matrix multiplication steps, making our reservoir relatively safe from high gain noise. To further reduce the impact, the amplified spontaneous noise (ASE) can be mitigated by the use of a band-pass filter [22] and in our simulations, we see good performance without requiring output probe powers over 31 dBm, which would result in a negligible noise level.
(a) Optical Lorenz System Test. NMSE=0.018
(b) Variation of Lorenz System Test NMSE with Pump Splitting Ratio
Figure 11: Lorenz-X Prediction Results to Demonstrate Accessibility of Negative Values
@@ IV.4 Noise Performance and Negative Values
Another source of noise is optical power detection noise. This is accounted for by a simple gaussian noise model. Figure 9 shows examples of the performance of the optical ESN with detector noise included, compared with Figures 3, 5, and 7. For this test, we assumed the noise property of $0.32\%$ uncertainty at 15501550nm wavelength, which is a standard treatment of considering detector noise [38]. In the presence of noise, the nonlinear capability is severely reduced as can be seen in the sine-square classification test results. This is mitigated by choosing an optimal $gL$ value with a large pump splitting ratio, as detailed in the previous section. We observe from Figure 10 that the average NMSE begins to trend upward, unlike the noise free cases of Figures 4, 6, and 8. The cause of this is likely that the saturation point for the specific task as a function of pump/input ratio is reached earlier because of the presence of detection noise.
@@
The optical ESN shows robustness in its time-series prediction ability, even in the presence of noise. The reported median NMSE in Figure 10 for the Mackey-Glass test is 0.15. This shows that the long-term memory capability is not diminished by the presence of noise, although there is a plateau in performance with varying pump splitting ratio, as can be seen from Figure 10 (c). We can see that overall, after accounting for power detection noise in the fiber setup, there is a reduction in performance of around 0.10.50.1-0.5 NMSE with more reduction in the classification task and less effect on prediction, implying that the efficiency of nonlinearity is affected more by the noise than the long and short term memory capabilities of the optical reservoir. These noise properties are for commercially available optical detectors. Hence, the negative impact of noise on performance can be further mitigated with the use of low noise detectors [39], [40].
One drawback of the selected benchmark tests is that they do not cover negative values. In order to demonstrate the ability of our design to access and predict negative inputs and targets, we use the Lorenz system test. It involves predicting the x-coordinate of a chaotic time series function [41]. The results for this test are shown in Figure 11.
We can see in Figure 11 that the architecture is able to produce negative outputs, and performs well for a task that involves negative values as well as positive values. We also see similar conclusions from the performance in this task as the other three: a gL value of $5-10W^{-1}$ is preferred and stronger pump splitting ratios provide better performance. This shows that negative values are also accessible for this architecture by using dBm units for the optical power.
V Conclusion
@@
We have provided the design for a feasible all-optical ESN with minimal resource overhead and energy requirements. For the nonlinear activation and matrix multiplication, we use SBS in two regimes: the large pump regime for linear behavior, and the self-pumped regime for nonlinear behavior. In the linear regime, tuning the pump power along with required attenuation using VOAs is enough to achieve the necessary multiplications. For the nonlinear activation step, according to reported achievable values of Brillouin gain in highly nonlinear fibers (HNLFs) [22], our design requires a fiber of length 363-6 m to achieve the required nonlinearity for good results, which is a very practical length of fiber. Such fibers also have a reasonable threshold power to show the required nonlinear output.
@@
Additionally, our design requires only a single Brillouin fiber amplifier as a nonlinear activation device, making it highly scalable. However, a trade-off in computation speed has been introduced to allow for this scalable design. It is possible to achieve faster operation speeds at the cost of scalability, but this also greatly increases the hardware resource requirement, so we have decided against this approach. The nonlinear activation is optimizable by tuning the $gL$ value of the SBS amplifier as well as the pump splitting ratio, which refers to the ratio of the pump to the input given to the self-pumped Brillouin amplifier. By doing so, we analyze the behavior of the optical ESN for multiple benchmark tests to find the most optimal activation function design, while ensuring low overall energy consumption.
@@
Our matrix multiplication and beam combination steps make use of standard fiber optical components such as VOAs, fiber beam combiners, 1:N splitters, combiners, and a 1:N switch. We see comparable results to a standard software ESN using a sigmoid activation function. To show this, we chose three tasks: classification of sine and square waves, prediction of the NARMA10 sequence, and prediction of the Mackey-Glass chaotic time series. Each benchmark task demonstrates the nonlinearity, polynomial algebra capability, and long-term memory of this ESN respectively, showing that universal computations are indeed accessible with this hardware. We validated our numerical simulations of the designed all-optical ESNs by comparing them to a software ESN with a sigmoid activation function. The results demonstrated that both systems achieve similar performance when using an equal number of nodes. This shows the feasibility and effectiveness of an all-optical approach to building an ESN.
VI Acknowledgment
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Biological and Environmental Research under Award Number DE-SC0025910.
@@ Appendix A SBS Transcendental Equation Derivation
$\int_{P_{\mathrm{s}}(0)}^{P_{\mathrm{s}}(z)}\frac{dP_{\mathrm{s}}}{P_{\mathrm{s}}(P_{\mathrm{s}}+P_{\mathrm{p}}-P_{\mathrm{out}})}=-\int_{0}^{z}gdz^{\prime}.$ (15)
$\implies ln\left(\frac{P_{\mathrm{s}}(z)(P_{\mathrm{in}}+P_{\mathrm{p}}-P_{\mathrm{out}})}{P_{\mathrm{in}}(P_{\mathrm{s}}(z)+P_{\mathrm{p}}-P_{\mathrm{out}})}\right)=-gCz.$ (16)
@@
$(P_{\mathrm{s}}(z)P_{\mathrm{in}}-P_{\mathrm{out}}P_{\mathrm{in}})e^{-(P_{\mathrm{p}}-P_{\mathrm{out}})gz}.$ (17)